OpenStudy (anonymous):

how do I factor with two variables a2 +7ab +10b2

OpenStudy (calculusfunctions):

Same as you would with one variable, except you have and b in your factors. a² + 7ab + 10b² = (a + 5b)(a + 2b) Do you get it?

OpenStudy (anonymous):

no can you break it down, where does the 5 come from, what is the answer?

OpenStudy (anonymous):

you have to split coefficient of ab into sum of 2 numbers so that their product becomes equal to the product of coefficients of b^2 and a^2

OpenStudy (calculusfunctions):

Check: (a + 5b)(a + 2b) = a(a + 2b) + 5b(a + 2b) = a² + 2ab + 5ab + 10b² = a² + 7ab + 10b²

OpenStudy (anonymous):

@kirkanthony see he has converted 7ab into 2ab+5ab because 2*5 = 1*10 :)

OpenStudy (calculusfunctions):

a² + 7ab + 10b² Since the coefficient of a² is 1. Find two integers which yield a product of 10 and a sum of 7. These two integers are clearly 2 and 5. correct? Now since the first term of the polynomial is a² and the last term is b², the first term of each binomial factor will be a and the second term of each binomial factor will contain a b.

OpenStudy (calculusfunctions):

Thanks @vikrantg4

OpenStudy (anonymous):

thank you

OpenStudy (calculusfunctions):

Hmmm...I don't get a medal?

OpenStudy (anonymous):

yes you do.